# MMGT: Motion Mask Guided Two-Stage Network for Co-Speech Gesture Video Generation

### v1

@@ Abstract
Co-Speech Gesture Video Generation aims to generate vivid speech videos from audio-driven still images, which is challenging due to the diversity of different parts of the body in terms of amplitude of motion, audio relevance, and detailed features. Relying solely on audio as the control signal often fails to capture large gesture movements in video, leading to more pronounced artifacts and distortions. Existing approaches typically address this issue by introducing additional a priori information, but this can limit the practical application of the task. Specifically, we propose a Motion Mask-Guided Two-Stage Network (MMGT) that uses audio, as well as motion masks and motion features generated from the audio signal to jointly drive the generation of synchronized speech gesture videos. In the first stage, the Spatial Mask-Guided Audio Pose Generation (SMGA) Network generates high-quality pose videos and motion masks from audio, effectively capturing large movements in key regions such as the face and gestures. In the second stage, we integrate the Motion Masked Hierarchical Audio Attention (MM-HAA) into the Stabilized Diffusion Video Generation model, overcoming limitations in fine-grained motion generation and region-specific detail control found in traditional methods. This guarantees high-quality, detailed upper-body video generation with accurate texture and motion details. Evaluations show improved video quality, lip-sync, and gesture. The model and code are available at (https://github.com/SIA-IDE/MMGT).
Index Terms:
Spatial Mask Guided Audio Pose Generation Network (SMGA), Unified model, Motion-masked hierarchical audio attention (MM-HAA)
Figure 1: Overview of existing models for generating videos of co-speech gestures. Compared with other methods [1, 2, 3], our method can generate videos of co-speech gestures for specified characters, and the generation for specific regions does not require additional prior information.
@@ I Introduction
[…]
[lines=2, findent=2mm, nindent=0mm]Co-speech face and gestures, as typical non-verbal behaviors, play an extremely important role in human-human communication[4, 5]. In a speech video, the combination of the speaker’s head movements and gestures can make a person’s speech more vivid and convey the meaning more clearly. Therefore, generating videos of the speaker’s gestures and head movements that are synchronized with the speech has attracted much attention from researchers. Unified audio-driven gesture and facial video generation has great potential for virtual worlds, Digital Human Development [6], and Multimedia Applications[7]. Although some current methods[8, 3, 1, 9, 2, 10, 11] can generate vivid speech gesture videos, unfortunately, videos generated solely using audio and static images often suffer from low quality[8, 11]. On the other hand, some methods[1, 3] achieve better quality but introduce additional a priori information beyond audio and inferred images. These a priori elements are not as easily accessible as audio, which limits the practical application of the technology.
@@
Existing methods mainly adopt end-to-end video generation frameworks. As shown in Fig. 1(a), MM-Diffusion [2] generates the full video directly from the audio based on a diffusion model, but does not introduce a reference image, resulting in an inability to effectively control the ID information of the characters in the video[11]. Diffted[8] have used audio multi-stage to generate gesture videos, and EchoMimicV2[1], CyberHost[3] have generated gesture videos by combining the audio with a priori information about the hand pose video or by utilizing codebook a priori enhancement techniques to generate gesture and face upper body videos. However, these methods still have several shortcomings. First, it is difficult to accurately control the synchronization of face and smooth movements of gestures simultaneously using only audio as a condition[2]. Second, most of the methods rely on other high-quality a priori conditions in addition to audio in the inference phase (see Fig.1(b)(c)), which is limited in practical applications. Third, if the model lacks region-specific constraints, as show in Fig. 1(a), the generated video tends to have large deformations and distortions in critical regions (e.g., face, gestures). For example, when there are large movements such as head rotations, the generated video may show head distortion [8].
@@
To address these challenges, we propose Motion Mask Guided Two-Stage Network (MMGT), which consists of two stages for video generation. In the first stage, we introduce the Spatial Mask Guided Audio Pose Generation Network (SMGA) to generate pose video guided by audio input. This pose video not only captures synchronized face but also naturally incorporates hand gestures. Based on this, as show in Algorithm 1, we further compute the maximum bounding box of each region to obtain motion mask.
@@
In the second stage, we propose a novel Motion Masked Hierarchical Audio Attention (MM-HAA) Module, which leverages audio to dynamically enhance the details of specific semantic regions while pose sequences drive the motion of static images. By introducing this module, our framework can more accurately control the generation of simultaneous interpretation videos, as show in Fig.2, significantly improving both realism and expressive fidelity.
@@
As shown in Fig. 1, comparison of input requirements and output capabilities across different methods. Unlike EchoMimicV2[1] and CyberHost[3], which require additional inputs such as hand pose sequences or body movement modules, our proposed method, MMGT, achieves synchronized generation of gestures and lip movements using only audio and a single reference image as inputs, without relying on any extra information. In summary, the main contributions of this work are as follows:
@@
1. We propose the Spatial Mask Guided Audio Pose Generation Network to generate a pose video and a motion mask from a speech signal and an initial pose. This unified design ensures natural and synchronized control of face and hand movements, resulting in cohesive and realistic co-speech animations.
@@
2. We introduce the Motion Masked Hierarchical Audio Attention Module for video generation. By integrating audio conditions with motion motion masks, the model adaptively focuses on specific semantic regions, such as hands, lips, and face, enabling fine-grained detail enhancement.
@@
3. Our network requires only audio and a single reference image as input to generate high-quality facial expression and gesture videos, without the need for additional prior information, significantly improving both practicality.
@@ II RELATED WORK
II-A Speech-driven Talking Face Video Generation
@@
In the field of speech-driven talking face video generation, early studies primarily focused on the lip region, given its strong correlation with audio [12]. These works leveraged GAN-based networks [13] to synchronize lip movements in the video with the corresponding audio [14]. Subsequently, researchers extended audio-lip synchronization to achieve synchronized generation of audio-driven head movements, face and emotions [15, 16, 17, 18]. Some studies leveraged pre-trained SD-diffusion models [19] to generate high-resolution and dynamic talking head videos from static images driven by audio [20, 21, 22, 16]. However, due to the limitations of existing models, it is challenging to extend face-specific video generation tasks to full upper-body video generation simply by retraining with a different dataset. Furthermore, current datasets also have inherent limitations: LRS2, VoxCeleb, and MEAD [23, 24, 25] only cover facial regions without including the full body, while the HDTF dataset [26] includes the entire upper body but lacks significant audio-correlated dynamic gestures. These issues collectively make it difficult for existing audio-driven talking face video generation methods to be adapted for gesture region synthesis. In contrast, our approach carefully considers the distinct correlations between facial and hand movements. By designing a well-structured network and utilizing the PATS dataset[27, 28, 29], our method effectively extends facial video generation to audio-driven upper-body video generation.
Figure 3: Overview of the Proposed MMGT Framework. The framework operates in two stages: In Stage I, the SMGA network generates motion feature videos, including the pose video $\hat{V}_{P}$ and motion mask $\hat{V}_{M}$, based on the input audio $a$ and initial pose $p^{(0)}$. In Stage II, the Denoising UNet utilizes $Z_{\text{pose}}$ and $Z_{\text{text}}$, while the ReferenceNet integrates $Z_{\text{pose}}$, $Z_{\text{text}}$, and $Z_{\text{img}}$ to produce the final predicted video $\hat{V}$. On the right, the MM-HAA module enhances $\hat{V}_{M}$ by aligning audio features $f_{a}$ with cross-attention embeddings $Z_{\text{CA}}$. The green double dashed line indicates the inference process, where $V_{P}$ and $V_{M}$ are replaced by $\hat{V}_{P}$ and $\hat{V}_{M}$ when transitioning from training to inference.
II-B Speech-driven Talking Gesture Generation
@@
In the field of gesture generation, early studies primarily relied on human skeleton data to generate vivid gesture sequences driven by audio [30, 31, 32]. Recent research has incorporated multiple modalities based on audio, such as emotion [33] and style [32], to improve the controllability of gesture sequence generation. Methods like EMAGE [34] and DiffSHEG [35] have extended co-speech gesture generation to include head movement synthesis. However, these approaches often disregard appearance information entirely [11], resulting in outputs that fail to align with users’ visual perception.
II-C Speech-driven Talking Gesture Video Generation
@@
Currently, gesture video generation methods can be broadly categorized into retrieval-based and generation-based methods. Zhou et al. [36] were the first to frame gesture video generation as a task of reproducing video frames aligned with audio, achieving high-quality gesture videos through a video frame retrieval method. TANGO [9] further advanced this approach by employing an action graph-based retrieval technique to generate co-speech body gesture videos. While retrieval-based methods produce satisfactory results with limited computational resources, they face notable limitations, including the inability to synthesize novel gestures and the lack of correlation modeling between the speaker’s face and audio.
@@
ANGIE[10] was the first to clearly define the problem of audio-driven gesture video generation, using unsupervised features MRAA[37] to simulate body movements. However, due to the linear nature of MRAA, it cannot represent complex regions, limiting the quality of video generation. Additionally, methods such as[11, 8] use non-linear TPS transformation[38] to decouple human ANGIE [10] was the first to explicitly define the task of audio-driven gesture video generation, leveraging the unsupervised MRAA framework [37] to simulate body movements. However, due to its linear nature, MRAA struggles to capture complex motion regions, thereby limiting the quality of generated videos. More recent methods, such as [11, 8], adopt non-linear TPS transformations [38] to decouple human motion keypoints from images. These methods generate full keypoint sequences driven by audio, which are then used to produce compact optical flow for animating images and creating smooth gesture videos. Despite their advancements, generation-based approaches face common challenges: they fail to simultaneously generate detailed finger and lip movements, exhibit significant artifacts and deformation, and suffer from limited generalizability.
Our method introduces the Spatial Mask Guided Audio Pose Generation Network (SMGA) network, meticulously designed to implicitly generate keypoint sequences from audio. By integrating a multimodal video generation framework enhanced with the Motion Masked Hierarchical Audio Attention (MM-HAA) mechanism, the proposed approach drives a single static image to simultaneously produce dynamic co-speech gestures and talking head videos. Beyond generating speech videos solely from audio, our method also supports pose-driven generation and joint driving with both pose and audio, providing greater flexibility and precise control over the output.
@@ III Proposed Method
We propose the MMGT network, as illustrated in Fig. 3, which ensures the synchronized control of both facial and hand movements, while also compensating for regional details. Given a segment of audio, we denote it as $\mathbf{a}=[a_{0},\dots,a_{N}]$ and a reference image $I_{0}\in\mathbb{R}^{3\times W\times H}$, our framework generates a co-speech video consisting of $N$ frames, denoted as $\hat{V}=[I_{0},\hat{I}_{1}\dots,\hat{I}_{N-1}]\in\mathbb{R}^{N\times 3\times W\times H}$. The co-speech video generation process is divided into two stages, motion feature generation (Stage I), motion-driven and detail refinement (Stage II). Therefore, our inference pipline can be formulated as
@@
$\hat{V}=\mathcal{G}_{v}(I_{0},\hat{V}_{P},\hat{V}_{M},a).$ (2)
where $E_{k}(\cdot)$ represents the initial pose extractor, responsible for obtaining the initial pose $p^{(0)}$ from the reference image $I_{0}$. In the first stage, our proposed $SMGA(\cdot)$ network, built upon the DiT architecture [39], takes the audio $a$ as a conditioning signal to generate the complete pose sequence $\hat{x}_{0}$. In the second stage, the video generation model $\mathcal{G}_{v}(\cdot)$ incorporates an reference-net and a temporal module based on the SD architecture [40]. It leverages the motion mask $\hat{V}_{M}$ generated in the first stage, along with the pose video and audio, as conditioning inputs to synthesize the final video $\hat{V}$.
III-A Overrall Framework Architecture
In the first stage, the SMGA network aims to capture the temporal and semantic relationships between motion features and audio input. As shown in Fig. 4, it consists of several key components. The motion block captures the correlation between audio and posture to obtain facial and overall posture features, and the merge block combines the two to obtain the final result.
In the second stage, the video generation network refines motion details using the generated motion features (motion mask, and pose videos), audio, and spatial cues from a reference image. This is achieved through a Denoising UNet, along with hierarchical attention module: Spatial Attention (SA) for texture details, Temporal Attention (TA) module for smooth transitions, and Motion Masked Hierarchical Audio Attention (MM-HAA) module for aligning audio with motion masks, ensuring temporal and spatial consistency in the generated video.
III-B Motion feature generation
As illustrated in Fig. 3, the primary objective of the initial phase is to synthesize pose videos and motion masks using audio inputs as the driving force. To achieve this, we have developed a Spatial Mask-Guided Audio Pose Generation (SMGA) Network, which generates both the complete pose video $\hat{V}_{P}$ and the corresponding motion mask $\hat{V}_{M}$. Among them, the pose video $\hat{V}_{P}$ effectively captures the basic dynamics of the motion. The motion mask $\hat{V}_{M}$ highlights key parts of significant motion, such as hands and faces, thus improving the accuracy of motion representation.
Figure 4: Architecture of the SMGA network.
Conditional Input Preparation. The initial pose $p^{(0)}$ is extracted from the reference image $I_{0}$ using a pre-trained keypoint detector[41] $E_{K}$,

### v2

@@ Abstract
Co-Speech Gesture Video Generation aims to generate vivid speech videos from audio-driven still images, which is challenging due to the diversity of body parts in terms of motion amplitude, audio relevance, and detailed features. Relying solely on audio as the control signal often fails to capture large gesture movements in videos, resulting in more noticeable artifacts and distortions. Existing approaches typically address this issue by adding extra prior inputs, but this can limit the practical application of the task. Specifically, we propose a Motion Mask-Guided Two-Stage Network (MMGT) that uses audio, along with motion masks and pose videos generated from the audio signal, to jointly generate synchronized speech gesture videos. In the first stage, the Spatial Mask-Guided Audio2Pose Generation (SMGA) Network generates high-quality pose videos and motion masks from audio, effectively capturing large movements in key regions such as the face and gestures. In the second stage, we integrate Motion Masked Hierarchical Audio Attention (MM-HAA) into the Stabilized Diffusion Video Generation model, addressing limitations in fine-grained motion generation and region-specific detail control found in traditional methods. This ensures high-quality, detailed upper-body videos with accurate textures and motion. Evaluations demonstrate improvements in video quality, lip-sync, and hand gestures. The model and code are available at (https://github.com/SIA-IDE/MMGT).
Figure 1: Overview of existing co-speech gesture video generation models. In contrast to other methods [31, 43, 25, 54], our approach (MMGT) does not require additional a priori inputs in the inference process.
@@ I Introduction
Co-speech facial expression and gestures, as typical non-verbal behaviors, play a crucial role in human communication[16, 7]. In a speech video, combining the speaker’s head movements and gestures can make a person’s speech more vivid and help convey the meaning more clearly. Therefore, generating videos of the speaker’s gestures and head movements that are synchronized with the speech has attracted significant interest from researchers. Unified co-speech gesture and face video generation has great potential for virtual worlds, Digital Human Development[33], and multimedia applications[34]. Although some current methods [18, 25, 31, 26, 43, 28, 17] can generate vivid speech gesture videos, unfortunately, videos generated solely from audio and reference image often lack quality [18, 17]. On the other hand, some approaches [31, 25] achieve better quality but require additional prior information beyond audio and inferred images. These prior elements are not as easily accessible as audio, which limits the practical use of the technology.
@@
Existing methods mainly use end-to-end video generation frameworks. As shown in Fig. 1(c), MM-Diffusion [43] generates the entire video directly from audio using a diffusion model but does not include a reference image, which limits effective control over the ID information of the characters in the video[17]. Diffted[18] uses multi-stage audio processing to generate gesture videos, while EchoMimicV2[31] and CyberHost[25] generate gesture videos by combining audio with prior information about hand pose videos or by motion frames and body movement map enhancement techniques to generate gesture and face videos.
However, these methods still have several limitations. First, it is challenging to accurately synchronize facial movements and smooth gestures using only audio as a condition [43]. Second, most methods depend on additional high-quality prior inputs besides audio during inference (see Fig. 1(a)(b)), which limits practical application. Third, lacking region-specific constraints, as shown in Fig. 1(c), the generated videos tend to exhibit significant deformations and distortions in critical regions such as the face and gestures. For instance, large movements like head rotations can cause head distortions in the generated videos [18].
@@
To address these challenges, we propose the Motion Mask Guided Two-Stage Network (MMGT), which consists of two stages for video generation. In the first stage, we introduce the Spatial Mask Guided Audio2Pose Generation Network (SMGA) to generate a pose video guided by audio input. This pose videos not only captures synchronized facial expressions but also naturally incorporates hand gestures. Based on this, as shown in Algorithm 1, we further compute the maximum bounding box for each region to obtain the motion mask.
@@
In the second stage, we introduce a novel Motion Masked Hierarchical Audio Attention (MM-HAA) Module, which uses audio to dynamically enhance the details of specific body regions while pose sequences guide the motion of reference image. By adding this module, our framework can more precisely control the generation of simultaneous interpretation videos, as shown in Fig. 2, significantly boosting both realism and expressive fidelity.
@@
As shown in Fig. 1, the comparison of input requirements and output capabilities across different methods is illustrated. Unlike EchoMimicV2[31] and CyberHost[25], which require additional inputs such as hand pose sequences or body movement modules, our proposed method, MMGT, achieves synchronized generation of gestures and lip movements using only audio and a single reference image as inputs, without relying on any extra prior information. In summary, the main contributions of this work are as follows:
@@
1. We propose the Spatial Mask Guided Audio2Pose Generation Network to generate a pose video and a motion mask from a speech signal and an initial pose. This integrated design provides natural and synchronized control of face and hand movements, leading to cohesive and realistic co-speech animations.
@@
2. We introduce the Motion Masked Hierarchical Audio Attention Module for video generation. By integrating audio conditions with motion masks, the model adaptively focuses on specific body regions, such as hand gestures, lips, and face, enabling fine-grained detail enhancement.
@@
3. Our MMGT network only needs audio and a single reference image to generate high-quality face and gesture videos, without extra prior inputs, significantly improving practicality.
@@
Figure 3: Overview of the Proposed MMGT Framework. The framework operates in two stages: In Stage I, the SMGA network generates motion feature (pose video $\hat{V}_{P}$, motion mask $\hat{V}_{M}$), based on the input audio $a$ and initial pose $p^{(0)}$. In Stage II, the Denoising UNet uses $Z_{\text{pose}}$ and $Z_{\text{text}}$, while ReferenceNet incorporates $Z_{\text{pose}}$, $Z_{\text{text}}$, and $Z_{\text{img}}$ to produce the final predicted video $\hat{V}$. On the right, the MM-HAA module enhances $\hat{V}_{M}$ by aligning audio features $f_{a}$ with cross-attention embeddings $Z_{\text{CA}}$. The green double-dashed line indicates the inference process, where $V_{P}$ and $V_{M}$ replace $\hat{V}_{P}$ and $\hat{V}_{M}$ when transitioning from training to inference.
@@ II RELATED WORK
II-A Co-speech Face Video Generation
In the field of co-speech face video generation, early studies [24, 37, 21] mainly concentrated on the lip region due to its strong link to audio. These works [37, 21] leveraged GAN-based networks [15] to synchronize lip movements in the video with the corresponding audio. Later, researchers expanded audio-lip synchronization to include synchronized generation of co-speech head movements, facial expressions, and emotions [59, 46, 30, 44]. Some studies [53, 54, 11, 46] utilized pre-trained SD-diffusion models [41] to generate high-resolution and dynamic talking head videos from reference image guided by audio.
@@ II-A Co-speech Face Video Generation
However, due to the limitations of current models, it is difficult to expand face-specific video generation tasks to full upper-body video creation simply by retraining with a different dataset, as shown in Fig. 1(d). These methods typically require the introduction of additional motion control features, otherwise, full-body control cannot be achieved. Furthermore, current datasets also have inherent limitations: LRS2, VoxCeleb, and MEAD [1, 32, 50] only cover face regions without including the full body, while the HDTF dataset [61] includes the entire upper body but lacks significant audio-correlated dynamic gestures. These issues collectively hinder the adaptation of existing co-speech face video generation methods for gesture region synthesis. By designing a well-structured network and utilizing the PATS dataset[3, 2, 4], our method effectively extends face video generation to co-speech upper-body video generation.
Figure 4: Comparison of motion masks (MMGT) and body movement maps CyberHost[25].
II-B Co-speech Gesture Generation
@@
In the field of gesture generation, early studies mainly relied on human skeleton data to generate lively gesture sequences driven by audio [58, 51, 55]. Recent research has incorporated multiple modalities based on audio, such as emotion [12] and style [55], to enhance the controllability of gesture sequence generation. Approaches like EMAGE [27] and DiffSHEG [9] have expanded co-speech gesture generation to include head movement synthesis. However, these methods often ignore appearance information entirely [17], leading to outputs that do not match users’ visual perception.
II-C Co-speech Gesture Video Generation
@@
Currently, gesture video generation methods are mainly divided into retrieval-based and generation-based methods. Zhou et al. [62] were the first to frame gesture video generation as a task of reproducing video frames aligned with audio, generating high-quality gesture videos through a video frame retrieval method. TANGO [26] further advanced this approach by employing an action graph-based retrieval technique to generate co-speech body gesture videos. Although retrieval-based methods deliver good results with limited computational resources, they have significant limitations, such as the inability to generate new gestures and the lack of modeling the connection between the speaker’s face and audio.
@@
ANGIE[28] was the first to clearly define the problem of co-speech gesture video generation, using unsupervised features MRAA[45] to simulate body movements. However, due to its linear nature, MRAA struggles to capture complex motion regions, thereby limiting the quality of generated videos. More recent methods, such as [17, 18], adopt non-linear TPS transformations [6] to decouple human motion keypoints from images. These methods generate full keypoint sequences driven by audio, which are then used to produce compact optical flow for animating images and creating smooth gesture videos.
Despite their advancements, generation-based approaches face common challenges: they fail to simultaneously generate detailed finger and lip movements, exhibit significant artifacts and deformation, and suffer from limited generalizability. CyberHost’s [25] performance gains largely stem from the integration of extra a priori cues during both training and inference, namely the Body Movement Map and Motion Frames shown in Fig. 1(a) and (b). EchoMimicV2 [31], meanwhile, synthesises gesture videos by conditioning on the input audio and hand-pose videos, then applies dedicated enhancement modules to generate co-speech videos. As shown in Fig. 4, CyberHost’s [25] input body movements maps are built from ground videos, which can only capture large-scale body movements. In contrast, our motion masks are derived from audio, which both accurately represent motion and enrich the detail of the corresponding regions.
@@ III Proposed Method
We propose the MMGT network, as shown in Fig. 3, which ensures synchronized control of both face and hand gestures while also compensating for regional details. Given a segment of audio denoted it as $\mathbf{a}=[a_{0},\dots,a_{N}]$ and a reference image $I_{0}\in\mathbb{R}^{3\times W\times H}$, our framework generates a co-speech video with $N$ frames, called $\hat{V}=[I_{0},\hat{I}_{1}\dots,\hat{I}_{N-1}]\in\mathbb{R}^{N\times 3\times W\times H}$. The process of co-speech video generation is divided into two stages: motion mask and pose videos generation (Stage I) and motion-driven and detail refinement (Stage II). Therefore, our inference pipeline can be formulated as
@@
$\hat{V}=\mathcal{G}_{v}(I_{0},\hat{V}_{P},\hat{V}_{M},a),$ (2)
where $E_{k}(\cdot)$ represents the initial pose extractor, responsible for obtaining the initial pose $p^{(0)}$ from the reference image $I_{0}$. In the first stage, our proposed $SMGA(\cdot)$ network, built upon the DiT architecture [35], takes audio $a$ as a conditioning signal to generate the complete pose sequence $\hat{x}_{0}$. In the second stage, the video generation model $\mathcal{G}_{v}(\cdot)$ incorporates an reference-net and a temporal module based on the SD architecture [42]. It leverages the motion mask $\hat{V}_{M}$ generated in the first stage, along with the pose video $\hat{V}_{P}$ and audio a\ a, as conditioning inputs to synthesize the final video $\hat{V}$.
III-A Overall Framework Architecture
In the first stage, the SMGA network aims to capture the temporal and semantic relationships between the pose sequence and the audio input. As shown in Fig. 5, it includes several key components. The motion block captures the correlation between audio and posture to obtain face and overall body features, and the merge block combines the two to generate the final output.
In the second stage, our framework leverages a Denoising UNet with three attention modules to refine motion details by integrating motion masks, pose videos, audio, and spatial cues from a reference image. Among these modules, the Spatial Attention (SA) and Temporal Attention (TA) are directly adopted from Animate-Anyone[20] to enhance texture details and ensure smooth transitions. Our key contribution lies in the Motion Masked Hierarchical Audio Attention (MM-HAA) module, which is inspired by the Hierarchical Audio Attention in Hallo [54] but innovatively replaces the static global mask with a dynamic motion mask derived from the pose sequence, as shown in Fig. 11. This design enables audio features to focus on actively moving regions in each frame, leading to more precise audiovisual alignment and improved temporal-spatial consistency in the generated video.
III-B Pose Video and Motion Mask Generation
As shown in Fig. 3, the main goal of the first stage is to synthesize pose videos and motion masks driven by audio inputs. To do this, we developed a Spatial Mask-Guided Audio2Pose Generation (SMGA) Network, which generates both the complete pose video $\hat{V}_{P}$ and the corresponding motion mask $\hat{V}_{M}$. Among them, the pose video $\hat{V}_{P}$ effectively captures the core dynamics of motion, while the motion mask $\hat{V}_{M}$ emphasizes key parts of significant movement, such as hand gestures and face, thereby enhancing the accuracy of motion representation.
Figure 5: Architecture of the $D_{SMGA}$.
Spatial Mask-guided Pose Video Generation. The initial pose $p^{(0)}$ is extracted from the reference image $I_{0}$ using a pre-trained keypoint detector[56] $E_{K}$,